# Adapting Image-to-Video Diffusion Models for Large-Motion Frame Interpolation

### v2

@@ I Introduction
Training a video diffusion model for video frame interpolation is a resource-intensive process that demands significant computational power to capture and model complex temporal dynamics between frames. To address this issue, we propose leveraging pre-trained image-to-video diffusion models and adapting them specifically for video frame interpolation tasks. Our approach is similar to Parameter-Efficient Fine-Tuning (PEFT)[11] techniques, enabling effective optimization of the base model while minimizing the need for extensive retraining.
@@ III-B Conditional Encoder
We present a novel plug-and-play conditional encoder module designed specifically for video frame interpolation shown in Fig 1. Firstly, we extract the first and last frames from the video data $\{I_{1},I_{2},\dots,I_{N}\}$, which results in $\{I_{1},I_{N}\}$. We design spatial and temporal branches to extract spatial-temporal features, aiming to enhance interpolation performance shown in Fig. 2. Additionally, a $1\times 1$ convolutional layer is employed to integrate conditional image CLIP [27] features for further processing.
Spatial Branch: Zero images are inserted between the first and last frames to represent the unconditional phase, with a concatenated binary mask to distinguish phases. The binary mask is set as $m=1$ for the conditional phase in the first and last frames, and $m=0$ for the unconditional phase in intermediate frames [28]. These components are concatenated along the channel dimension. A pre-trained autoencoder is then utilized to encode the combined inputs, extracting spatial features.
@@ IV-C Qualitative Result
TABLE II: Ablation study of proposed components, evaluating the middle of 5 generated frames at $512\times 512$ resolution.
PSNR↑ SSIM↑ LPIPS↓ FID↓ FVD↓
Baseline 19.02 0.5145 0.3161 16.50 134.6
Ours 20.28 0.5375 0.3069 16.27 118.05
@@ IV-D Ablation study
We conducted experiments on the proposed components of our method to perform an ablation study. Refer to Table II for detailed results. For a fair comparison of the proposed method, we discard the $256\times 256$ versions of the DAVIS-7 and UCF101-7 datasets. Instead, we re-crop the DAVIS 2017 dataset to $512\times 512$ resolution, retaining sequences of 7 frames, where the model is tasked with predicting the middle 5 frames. The baseline model includes only the spatial branch, enabling an image-to-video model to adapt to video frame interpolation. In contrast, our method integrates the temporal branch and the cross-frame attention module. Experimental results demonstrate that the proposed components enhance the performance of the baseline model.

### v3

@@ I Introduction
Training a video diffusion model for video frame interpolation is a resource-intensive process that demands significant computational power to capture and model complex temporal dynamics between frames. To address this issue, we propose leveraging pre-trained image-to-video diffusion models and adapting them specifically for video frame interpolation tasks. Our approach is similar to Parameter-Efficient Fine-Tuning (PEFT) techniques, enabling effective optimization of the base model while minimizing the need for extensive retraining.
@@ III-B Conditional Encoder
We present a novel plug-and-play conditional encoder module designed specifically for video frame interpolation, As discussed in previous work[26], adding noised latents to the conditional encoder can cause the model to overlook these samples, leading to inconsistencies in frame interpolation during training. Therefore, we also eliminate the noised latents and instead rely solely on the conditional inputs as shown in Fig 1. Firstly, we extract the first and last frames from the video data $\{I_{1},I_{2},\dots,I_{N}\}$, which results in $\{I_{1},I_{N}\}$. We design spatial and temporal branches to extract spatial-temporal features, aiming to enhance interpolation performance shown in Fig. 2. Additionally, a $1\times 1$ convolutional layer is employed to integrate conditional image CLIP [27] features for further processing.
Spatial Branch: Zero images are inserted between the first and last frames to represent the unconditional phase, with a concatenated binary mask to distinguish phases. The binary mask is set as $m=1$ for the conditional phase in the first and last frames, and $m=0$ for the unconditional phase in intermediate frames. These components are concatenated along the channel dimension. A pre-trained autoencoder is then utilized to encode the combined inputs, extracting spatial features.
@@ IV-B Quantitative Result
TABLE II: Ablation study of proposed components, evaluating the middle of 7 generated frames at 512 × 512 resolution
PSNR↑ SSIM↑ LPIPS↓ FID↓ FVD↓
Baseline 19.01 0.4839 0.3502 24.52 183.1
w/ Cross-frame attention 18.86 0.4794 0.3541 24.35 157.3
w/ Temporal features 19.03 0.4869 0.3520 24.69 169.4
Ours 19.18 0.4874 0.3366 22.34 148.2
@@ IV-D Ablation study
We conducted experiments on the proposed components of our method to perform an ablation study. Detailed results are provided in Table II. To mitigate the impact of down sampling on model performance, we discard the $256\times 256$ version of DAVIS-7. Instead, we re-crop the DAVIS 2017 dataset to a resolution of $512\times 512$, retaining sequences of 9 frames, where the model is tasked with predicting the middle 7 frames. The baseline model includes only the spatial branch, enabling an image-to-video model to adapt to video frame interpolation. In contrast, our method integrates the temporal branch and the cross-frame attention module. Experimental results demonstrate that the proposed components significantly enhance the performance of the baseline model.